How many grams of potassium nitrate and water must be used to prepare 175 g of a 22.5% solution
1
Expert's answer
2020-04-09T13:11:48-0400
"m(KNO_3) = w*m_{solution}"

"m(KNO_3) = 0.225*175 = 39.375 g"